OpenStudy (anonymous):

Solve each quadratic equation using factoring?? Pleasee help :) x^2+8x+16=25

OpenStudy (cgreenwade2000):

ax^2+bx+c=0

OpenStudy (anonymous):

First subtract 25 from both sides

OpenStudy (anonymous):

okay then....

OpenStudy (cgreenwade2000):

I see what you're doing. You want us to solve them for you. You could have learned how to do it by now with all the questions you're asking us to solve.

OpenStudy (anonymous):

xx^2 + 8x -9 = 0 Factor it (x + ?)(x - ?) Multiplies to make -9 and subtracts to give you 8

OpenStudy (anonymous):

no i'm actually doing the problem step my step like he/she is telling me soooooo

OpenStudy (anonymous):

So what are the factors?

OpenStudy (anonymous):

@lisapeajac why is there two x's at the beginnning?

OpenStudy (anonymous):

cause i can't type

OpenStudy (anonymous):

so the factors will be 1,9

OpenStudy (anonymous):

Yep. Which is positive and which is negative?

OpenStudy (anonymous):

the 1 is postive.. the 9 is negative

OpenStudy (cgreenwade2000):

Close. The 1 is negative and the 9 is positive because 9+(-1) = 8

OpenStudy (anonymous):

It factors like this: (x + 9)( x - 1)=0 and then when you solve it, the solutions are -9 and 1.

OpenStudy (anonymous):

thanks you both

OpenStudy (anonymous):

You're welcome